Two spherical objects have equal masses and experience a gravitational force of 25 N towards one another. Their centers are 36 cm apart. Determine each of their masses.


1
Expert's answer
2021-12-06T09:43:47-0500

"F=Gmm/r^2\to m=\sqrt{F\cdot r^2/G}=\sqrt{25\cdot 0.36^2/(6.67\cdot10^{-11})}\approx220400\ (kg)"